# Evaluating Decision Optimality of Autonomous Driving via Metamorphic Testing

### v1

@@ 7. Conclusion
In this paper, we propose the first study to detect non-optimal decisions in ADSs, which lacks the oracles to verify the optimality of the generated paths. To address this issue, we propose a metamorphic testing method, Decictor, incorporating a specifically designed metamorphic relation (MR). Decictor comprises three main components: Non-invasive mutation, which generates new scenarios satisfying the constraints described in the MR; MR Check component, verifying whether the scenario’s execution adheres to the MR; Feedback-guided selection component, responsible for selecting scenarios for the subsequent mutation. The experimental results demonstrate the effectiveness and efficiency of Decictor, as well as the usefulness of each component in Decictor.

### v3

@@ VII Conclusion
In this paper, we propose the first study to evaluate the robustness of ADSs’ path-planning decisions. To address this issue, we develop a testing tool Decictor, which comprises three main components: Non-invasive Mutation, generating new scenarios that are unlikely to affect the original optimal paths; Consistency Check, determining whether the driving paths in the mutated scenarios are consistent with the original optimal one; Feedback-guided Selection, responsible for selecting better scenarios for the subsequent mutation. The experimental results demonstrate the effectiveness and efficiency of Decictor, as well as the usefulness of each component in Decictor.
Acknowledgment
This research is partially supported by the Ministry of Education, Singapore under its Academic Research Fund Tier 2 (Proposal ID: T2EP20223-0043), the Lee Kong Chian Fellowship, the National Natural Science Foundation of China (62102283) and the Beijing Nova Program. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the Ministry of Education, Singapore.